i have the problem:
ln(x+1)^2=2
i worked it out as:
ln(x+1)^2=2
(x+1)^2=e^2
2(x+1)=7.38906
2x+2=7.38906
2x\2 = 5.38906
x=2.69453
i know i did something wrong becausemy answer is not working out.
help me plz!

Answer by Alan3354(21583)

ln(x+1)^2=2
i worked it out as:
ln(x+1)^2=2
(x+1)^2=e^2 You changed from exponent 2 to multiply by 2
-----------
2(x+1)=7.38906
2x+2=7.38906
----------
----------
ln(x+1)^2=2
2ln(x+1) = 2
ln(x+1) = 1
x+1 = e
x = e-1